# Experimental practical quantum tokens with transaction time advantage

### v1

@@ Abstract
Laboratory of Radiation Detection and Medical Imaging and School of Health Science and Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
State Key Laboratory of Functional Materials for Informatics, Shanghai Institute of Microsystem and Information Technology, Chinese Academy of Sciences, Shanghai 200050, China
@@
Here we present a full experimental implementation of the quantum S-token scheme of Refs. [11, 13, 14], demonstrating for the first time quantum tokens with near-perfect security and user privacy and near-instant validation in a realistic photonic setup that considers losses, errors and experimental imperfections. We present a new security analysis based on maximal confidence quantum measurement bounds [29] to prove unforgeability. Due to a high detection heralding efficiency of 88.24%, our implementation allows the correct presentation and validation of tokens while guaranteeing user privacy by not requiring the user to report losses when the token is generated. This prevents multi-photon attacks [15], to which many previous experimental demonstrations of mistrustful quantum cryptography (e.g., [30, 31, 32, 33, 34, 14, 35]) were vulnerable. Moreover, our experiment demonstrates for the first time a secure quantum token scheme that achieves faster validation than classical cross-checking can, given relativistic signalling constraints.
@@
By developing a data transmission and processing board with a 10 Gigabaud rate, we demonstrate quantum token transactions using a metropolitan fiber network, showing transaction time advantages over any classical cross-checking token schemes. We also demonstrate these advantages in short- range urban fiber optic networks.
@@
The scheme.— Following Refs. [11, 13, 14], Bob (the bank) sends Alice (the client) the quantum token via a short-range quantum channel. Alice measures the quantum states immediately upon reception without using quantum memories. The scheme proceeds with classical communications, without further quantum communications.
@@
We consider a two-node network scenario (see Fig.  S3 in SI) involving small spatial regions $L_{0},L_{1}$, defined in an agreed reference frame $F$. Alice (Bob) comprises two collaborating and mutually trusting agents or laboratories $A_{0}$ and $A_{1}$ ($B_{0}$ and $B_{1}$), located in $L_{0}$ and $L_{1}$ respectively connected via secure and authenticated classical channels (implemented via pre-distributed secret keys, for instance). In general, Alice and Bob may be companies or governments with many distributed trusted agents: the two-node two-agent scenario models the simplest case (see SI for extensions to larger networks). Note that any money or token scheme applicable in time-critical scenarios on extended networks needs distributed trusted agents, since tokens can be transmitted at near light speed while individual agents cannot. We define the space-time region $R_{i}$ to comprise the location $L_{i}$ within a time interval $\Delta T$ beginning at a time $T$ in $F$. Let $\Delta T_{\text{comm}}$ be the time taken for A0 to communicate a bit to A1. The values of $\Delta T$ and $\Delta T_{\text{comm}}$ are agreed in advance by Alice and Bob, with $\Delta T$ large enough to allow Alice to present the token to Bob. We assume that $T$ is only communicated by Alice to Bob at step 3 in Table 1 via the relation
@@
    Below, $\textbf{x},\textbf{x}^{\prime},\textbf{x}^{0},\textbf{x}^{1},\textbf{t},\textbf{u}$ are $N-$bit strings, $z,b,c,d_{0},d_{1}$ are bits, and $s_{k}$ is the $k$th bit of the string s.
@@
We employed the only known perfect protection against general multiphoton attacks [15], with A0 not reporting losses and accepting all pulses transmitted by B0. A0 assigned random measurement outcomes $x_{i}\in\{0,1\}$ for the pulses activating none or both of her detectors, introducing errors for these cases with probability close to $50\%$. Following Ref [38] and employing improved experimental techniques, the system efficiency of the heralded single-photon source is $88.24\%$ utilizing SNSPDs with over $95\%$ efficiency. This high efficiency gives a 7 standard deviation bound for the overall error rate of only $E=6.2550\%$. By setting $\gamma_{\text{err}}=9.4\%,$ we guarantee our implementation to be $2.1\times 10^{-11}$-correct (Bob rejects a valid token with a probability ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.1}\times 10^{-11}$). Our implementation is proved $\beta_{E}$-private (Bob learns Alice’s chosen presentation region before she presents with a probability ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}\frac{1}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta_{E}}$), given the perfect protection against multiphoton attacks. Here the bias $\beta_{\text{E}}=10^{-5}$ of the bit $z$ encoding A0’s measurement basis in our implementation and can be made arbitrarily small by pre-processing. This assumes that B0 cannot exploit side-channels to obtain information about A0’s measurement basis nor implement clock synchronization attacks to obtain information about A0’s chosen presentation location prematurely. Seven standard deviation upper bounds on B0’s biases in selecting the preparation basis and state and the proportion of multi-photon heralded pulses were respectively $\beta_{\text{PB}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.001360}$, $\beta_{\text{PS}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.001120}$, $P_{\text{noqub}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4.9\times 10^{-5}}$.
@@
The uncertainty angle in Bob’s state preparation is guaranteed to be $\leq\theta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}5.115515^{\circ}}$ with a probability $\geq(1-P_{\theta})$, where $P_{\theta}=0.027$. This shows our scheme to be $5.52\times 10^{-9}$-unforgeable (Alice succeeds in getting Bob to validate tokens at both presentation regions with a probability $\leq 5.52\times 10^{-9}$). This follows from a novel security analysis based on bounds on Alice’s maximum confidence quantum measurement [29] for each pulse ( see Methods).
@@
The classical phase (steps 3-6) was implemented with high-speed electronic boards and communication links within optical fibre networks. To demonstrate the time advantage over classical cross-checking protocols, each Ai communicates to Bi at 10 Gbps and the Bi perform real-time local validation using high-speed field-programmable gate arrays (FPGAs). The total duration of the classical processing without considering the communication time between $L_{0}$ and $L_{1}$ is $\Delta T_{\text{proc}}\approx 1.5~{}\mu\text{s}$.
The relationship between transaction time and distance is illustrated in Fig. 3(a) and (b). With straight fiber optic channels, quantum and comparative advantage can be demonstrated at approximately 0.3 km and 0.9 km, respectively (see Methods). Even though real optical fiber channels are not straight, our experiment demonstrates quantum and comparative advantage respectively in an intra- and inter-city network.
@@
We also demonstrated comparative advantage between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China, separated by 51.60 km and connected by a 60.54 km field-deployable optical fiber, as shown in Fig. 2(b). We ran the transactions 20 times , with all error rates below $\gamma_{err}=9.4\%$   for the tokens x. The respective average error rate and transaction time were $6.00\%$ and $304.202\pm 0.011~{}\mu s$, achieving the comparative advantage of $39.798\pm 0.011~{}\mu s$, shown in Fig. 3(b).
@@
Our theoretical and experimental techniques apply more broadly to mistrustful quantum cryptography. To our knowledge, this is the first mistrustful quantum cryptography experiment perfectly closing the multiphoton attacks loophole. We have advanced the theory of quantum token security [14] by new results using bounds on Alice’s maximum confidence quantum measurement [29] for each pulse, which also apply to other quantum token (e.g., [26]) and mistrustful quantum cryptographic schemes (e.g.,[31, 39, 40]). These bounds imply near-perfect token unforgeability allowing for experimentally quantified error types and loss levels. Unlike previous analyses (e.g., [14, 31]), they do not assume that source qubit states belong to orthonormal bases. We characterized the deviation from BB84 states and used this in our security proof, going substantially beyond previous security analyses (e.g., [41, 42, 43, 31, 32, 34, 26, 14, 35]) in mistrustful quantum cryptography. To our knowledge, no previous security analysis has allowed for general deviations from the set of states stipulated by an ideal protocol, measured these deviations experimentally, and based security bounds on these empirical data; without these results, claimed security bounds are not reliable. (See Methods and SI.) Our experiment is also the first demonstration of a quantitative quantum time advantage in relativistic cryptography, showing the enhanced cryptographic power of simultaneously considering quantum and relativistic physics.
@@
Acknowledgement.—This work was supported by the National Key Research and Development Program of China (2020YFA0309704), the Innovation Program for Quantum Science and Technology (2021ZD0300800), the National Natural Science Foundation of China (12374470), the Key R&D Plan of Shandong Province (2021ZDPT01), the Key Area R&D Program of Guangdong Province (2020B0303010001), Shandong Provincial Natural Science Foundation (ZR2021LLZ005), Y.-F.J and Q.Z acknowledge support from the Taishan Scholars Program. A.K. and D.P.-G. acknowledge support from the UK Quantum Communications Hub grant no. EP/T001011/1. G.C. was supported by a studentship from the Engineering and Physical Sciences Research Council. A.K. was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science.
@@
Figure 1: Diagram of the quantum token generation using a heralded single-photon source. A distributed feedback (DFB) laser with a central wavelength of 1560 nm is used as the pump. The laser emits a pulse with a width of 5 ns and a repetition rate of 500 KHz. The pump is frequency-doubled with a periodically poled MgO-doped lithium niobate (PPMgLN) crystal and filtered by a wavelength division multiplex (WDM) to create the 780 nm pump. The photon pairs are generated through the spontaneous parametric down-conversion (SPDC) process, utilizing the Type-II periodically poled potassium titanyl phosphate (PPKTP) crystal. One photon is used to trigger the superconducting nanowire single-photon detector (SNSPD), while the other is sent to Alice. Alice measures the received qubits using a half-wave plate (HWP) and a polarizing beam splitter (PBS), and the photons are detected by SNSPDs.
@@
Figure 2: Field deployment of the S-token. (a) A satellite image shows the S-token setup in the fiber optic network within the city of Jinan, Shandong Province, China. The fibre length covers 2,766 m, whereas the corresponding direct free space distance is around 425 m. (b) A satellite image displays the S-token setup with field-deployed fibre between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China. The fiber length is 60.54 km, while the direct free space distance between them is about 51.60 km.
@@
Figure 3: The results of the time advantage. (a) The black and red lines represent the transaction time based on classical cross-checking and S-token in fibre, respectively. The squares are the results of the transaction in Jinan with the fibre length of 2.766 km. (b) the black line represents the transaction time based on classical cross-checking in free space. The red line is the transaction time based on the S-token in fibre. In the inserted figure, the comparative advantage can be seen to be approximately 0.9 km in the ideal scenario. The stars are the experimental results of the completion between Yiyuan and Mazhan. It is worth noting that the optical fibre length (60.54 km) is longer than the free-space distance (51.60 km).

### v2

@@ Abstract
Quantum Medical Sensing Laboratory and School of Health Science and Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
Shanghai Key Laboratory of Superconductor Integrated Circuit Technology, Shanghai Institute of Microsystem and Information Technology, Chinese Academy of Sciences, Shanghai 200050, China
@@
Here we present a full experimental implementation of the quantum S-token scheme of Refs. [11, 13, 14], demonstrating for the first time quantum tokens with near-perfect security and user privacy and near-instant validation in a realistic photonic setup that considers losses, errors and experimental imperfections. We present a new security analysis based on maximal confidence quantum measurement bounds [29] to prove unforgeability. Due to a high detection heralding efficiency of 88.24%, our implementation allows the correct presentation and validation of tokens while guaranteeing user privacy by not requiring the user to report losses when the token is generated. This prevents multi-photon attacks [15], to which many previous experimental demonstrations of mistrustful quantum cryptography (e.g., [30, 31, 32, 33, 34, 14]) were vulnerable. Moreover, our experiment demonstrates for the first time a secure quantum token scheme that achieves faster validation than classical cross-checking can, given relativistic signalling constraints.
@@
By developing a data transmission and processing board with a 10 Gigabaud rate, we demonstrate quantum token transactions using a metropolitan fibre network, showing transaction time advantages over any classical cross-checking token schemes. We also demonstrate these advantages in short- range urban fibre optic networks.
@@
The scheme.— Following Refs. [11, 13, 14], Bob (the bank) sends Alice (the client) the quantum token via a short-range quantum channel. Alice measures the quantum states immediately upon reception without using quantum memories. The scheme then proceeds with classical communications, without further quantum communications.
@@
We consider a two-node network scenario (see Fig.  S3 in SI) involving small spatial regions $L_{0},L_{1}$, defined in an agreed reference frame $F$. Alice (Bob) comprises two collaborating and mutually trusting agents or laboratories A0 and A1 (B0 and B1), located in $L_{0}$ and $L_{1}$ respectively connected via secure and authenticated classical channels (implemented via pre-distributed secret keys, for instance). In general, Alice and Bob may be companies or governments with many distributed trusted agents: the two-node two-agent scenario models the simplest case (see SI for extensions to larger networks). Note that any money or token scheme applicable in time-critical scenarios on extended networks needs distributed trusted agents, since tokens can be transmitted at near light speed while individual agents cannot. We define the space-time region $R_{i}$ to comprise the location $L_{i}$ within a time interval $\Delta T$ beginning at a time $T$ in $F$. Let $\Delta T_{\text{comm}}$ be the time taken for A0 to communicate a bit to A1. The values of $\Delta T$ and $\Delta T_{\text{comm}}$ are agreed in advance by Alice and Bob, with $\Delta T$ large enough to allow Alice to present the token to Bob. We assume that $T$ is only communicated by Alice to Bob at step 3 in Table 1 via the relation
@@
    {spacing}1.0 Below, $\textbf{x},\textbf{x}^{\prime},\textbf{x}^{0},\textbf{x}^{1},\textbf{t},\textbf{u}$ are $N-$bit strings, $z,b,c,d_{0},d_{1}$ are bits, and $s_{k}$ is the $k$th bit of the string s.
@@
We employed the only known perfect protection against general multiphoton attacks [15], with A0 not reporting losses and accepting all pulses transmitted by B0. A0 assigned random measurement outcomes $x_{i}\in\{0,1\}$ for the pulses activating none or both of her detectors, introducing errors for these cases with probability close to $50\%$. Following Ref [37] and employing improved experimental techniques, we measured the final system efficiency of $88.24\%$, including the heralded single-photon source, the encoding by Pockels cells and the measurement by A0. This high system efficiency gives a 7 standard deviation bound for the overall error rate of only $E=6.2550\%$. By setting $\gamma_{\text{err}}=9.4\%$ and $N=10,048$ bits, we guarantee our implementation to be $2.1\times 10^{-11}$-correct (Bob rejects a valid token with a probability ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.1}\times 10^{-11}$). Our implementation is proved $\beta_{E}$-private (Bob learns Alice’s chosen presentation region before she presents with a probability ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq}\frac{1}{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta_{E}}$), given the perfect protection against multiphoton attacks. Here the bias $\beta_{\text{E}}=10^{-5}$ of the bit $z$ encoding A0’s measurement basis in our implementation and can be made arbitrarily small by pre-processing. This assumes that B0 cannot exploit side-channels to obtain information about A0’s measurement basis nor implement clock synchronization attacks to obtain information about A0’s chosen presentation location prematurely. Seven standard deviation upper bounds on B0’s biases in selecting the preparation basis and state and the proportion of multi-photon heralded pulses were respectively $\beta_{\text{PB}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.001360}$, $\beta_{\text{PS}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}0.001120}$, $P_{\text{noqub}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}4.9\times 10^{-5}}$.
@@
Based on the measurement, the uncertainty angle in Bob’s state preparation is guaranteed to be $\leq\theta={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}5.115515^{\circ}}$ with a probability $\geq(1-P_{\theta})$, where $P_{\theta}=0.027$ (see SI). This shows our scheme to be $5.52\times 10^{-9}$-unforgeable (Alice succeeds in getting Bob to validate tokens at both presentation regions with a probability $\leq 5.52\times 10^{-9}$). This follows from a novel security analysis based on bounds on Alice’s maximum confidence quantum measurement [29] for each pulse ( see Methods).
@@
The classical phase (steps 3-6) was implemented with optical fibre communication links and high-speed electronic boards. To demonstrate the time advantage over classical cross-checking protocols, each Ai communicates to Bi at 10 Gbps and the Bi perform real-time local validation using high-speed field-programmable gate arrays (FPGAs). The total duration of the classical processing for $N=10,048$ bits without considering the communication time between $L_{0}$ and $L_{1}$ is $\Delta T_{\text{proc}}\approx 1.5~{}\mu\text{s}$.
@@
We further demonstrated comparative advantage between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China, separated by 51.60 km and connected by a 60.54 km field-deployable optical fibre, as shown in Fig. 2(b). We ran the transactions 20 times and measured the average error rate and transaction time were $6.00\%$ and $304.20\pm 0.01~{}\mu s$, achieving the comparative advantage of $39.80\pm 0.01~{}\mu s$, shown in Fig. 3(b).
The relationship between transaction time and distance is illustrated in Fig. 3(a) and (b). Assume straight fibre optic channels, the quantum and comparative advantage can be achieved at approximately 0.3 km and 0.9 km, respectively (see Methods). Even though real optical fibre channels are not straight, our experiment demonstrates quantum and comparative advantage respectively in an intra- and inter-city network.
@@
Our theoretical and experimental techniques apply more broadly to mistrustful quantum cryptography. To our knowledge, this is the first mistrustful relativistic quantum cryptography experiment perfectly closing the multiphoton attacks loophole. We have advanced the theory of quantum token security [14] by new results using bounds on Alice’s maximum confidence quantum measurement [29] for each pulse, which also apply to other quantum token (e.g., [26]) and mistrustful quantum cryptographic schemes (e.g.,[31, 39, 40]). These bounds imply near-perfect token unforgeability allowing for experimentally quantified error types and loss levels. Unlike previous analyses (e.g., [14, 31]), they do not assume that source qubit states belong to orthonormal bases. We characterized the deviation from BB84 states and used this in our security proof, going substantially beyond previous security analyses (e.g., [41, 42, 43, 31, 32, 34, 26, 14, 38]) in mistrustful quantum cryptography. To our knowledge, no previous security analysis has allowed for general deviations from the set of states stipulated by an ideal protocol, measured these deviations experimentally, and based security bounds on these empirical data; without these results, claimed security bounds are not reliable. (See Methods and SI.) Our experiment is also the first demonstration of a quantitative quantum time advantage in relativistic cryptography, showing the enhanced cryptographic power of simultaneously considering quantum and relativistic physics.
@@
Acknowledgement This work was supported by the National Key Research and Development Program of China (2020YFA0309704), the Innovation Program for Quantum Science and Technology (2021ZD0300800, 2023ZD0300100), the National Natural Science Foundation of China (12374470), the Key R&D Plan of Shandong Province (2021ZDPT01), the Key Area R&D Program of Guangdong Province (2020B0303010001), Shandong Provincial Natural Science Foundation (ZR2021LLZ005), Y.-F.J and Q.Z acknowledge support from the Taishan Scholars Program. A.K. and D.P.-G. acknowledge support from the UK Quantum Communications Hub grant no. EP/T001011/1. G.C. was supported by a studentship from the Engineering and Physical Sciences Research Council. A.K. was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Research, Innovation and Science.
@@
Figure 1: Diagram of the quantum phase of the quantum token generation using a heralded single-photon source. A distributed feedback laser (DFB) with a central wavelength of 1560 nm is used as the pump, driven by a pulse generator (PG). The laser emits a pulse with a width of 5 ns and a repetition rate of 500 KHz. The pump is frequency-doubled with a periodically poled MgO-doped lithium niobate (PPLN) crystal and filtered by a wavelength division multiplex (WDM) to create the 780 nm pump. The photon pairs are generated through the spontaneous parametric down-conversion (SPDC) process, utilizing the Type-II periodically poled potassium titanyl phosphate (PPKTP) crystal. One photon is detected by the superconducting nanowire single-photon detector (SNSPD) to generate a trigger signal. The other is encoded with the basis $u_{i}$ and bit $t_{i}$ using the two Pockels cells respectively, which are driven by quantum random number generators (QRNGs). The XOR is an Exclusive OR operation implemented with electronic. A0 measures the received qubits using a half-wave plate (HWP) and a polarizing beam splitter (PBS), and the photons are detected by SNSPDs.
@@
Figure 2: Field deployment of the S-token. (a) A satellite image shows the S-token setup in the fibre optic network within the city of Jinan, Shandong Province, China, for demonstrating the quantum advantage. The fibre length is 2.766 km, whereas the corresponding direct free space distance is around 425 m. (b) A satellite image displays the S-token setup with field-deployed fibre between Yiyuan ($36^{\circ}10^{\prime}50.4^{\prime\prime}N,118^{\circ}12^{\prime}10^{\prime\prime}E$) and Mazhan ($36^{\circ}0^{\prime}19^{\prime\prime}N,118^{\circ}42^{\prime}35^{\prime\prime}E$) in Shandong Province of China, for demonstrating the comparative advantage. The fibre length is 60.54 km, while the direct free space distance between them is about 51.60 km.
@@
Figure 3: The results of the time advantage. (a) The red and black lines are the transaction time based on the S-token and classical cross-checking through fibre. The red square is the measured transaction time of the S-token through 2.766 km fibre link, the black square is the ideal transaction time of classical cross-checking via the same fibre length. (b) The red and black lines are the transaction time based on the S-token and classical cross-checking through fibre and free space channel, respectively. The red star is the measured transaction time of the S-token through 60.54 km fibre link, the black star is the ideal transaction time of classical cross-checking through 51.60 km free space link. Insert: the comparative advantage is achieved for distances longer than 0.9 km.